OpenStudy (anonymous):

An automobile is approaching you at a speed of 90.0 km/h and sounding its horn. The fundamental frequency of the horn sounds to you like 268 Hz. If the real fundamental frequency of the horn is 248 Hz, what is the speed of sound?

OpenStudy (anonymous):

my attempt: fd=(Vw/(Vw-Vs))*fs Vs=25.0m/s fd=268Hz fs=248Hz Vw=? i am not sure how to manipulate the equation

OpenStudy (aaronq):

\(\sf f_+=\dfrac{f_o}{1-\dfrac{v_s}{v}}\) \(\sf f_+*(1-\dfrac{v_s}{v})=f_o\) \(\sf (1-\dfrac{v_s}{v})=\dfrac{f_o}{f_+}\) \(\sf \dfrac{v_s}{v}=\dfrac{f_o}{f_+}+1\) \(\sf v_s=v*(\dfrac{f_o}{f_+}+1)\)

OpenStudy (anonymous):

?